# The Influence of Exclusion Zones on the Coexistence of Predator and Prey with an Allee Effect

### v1

@@ 1.5. Related literature
Specifically, our paper addresses the gaps in the literature through three principal advances. First, we establish the existence of coexistence equilibria without appealing to bifurcation from semi-trivial states. Using a novel topological-degree argument, we prove the existence of nontrivial steady states in parameter regimes that may be remote from classical bifurcation points. This yields a more global and geometrically robust understanding of coexistence, complementing and extending the local bifurcation analyses of, e.g., [13, 12, 36, 57]. Second, we develop a rigorous thin-limit analysis showing how coexistence persists and transforms as the predator region becomes vanishingly thin. This asymptotic framework connects the fully coupled PDE model to lower-dimensional effective equations, revealing new behaviors not captured by prior protection-zone studies. Finally, our numerical investigations uncover a variety of nonlinear phenomena including multiple coexistence branches, turning-point bifurcations, Hopf bifurcations, and transitions between extinction and persistence regimes. These phenomena highlight the rich dynamical landscape made possible by the coupling of Allee-type prey dynamics and spatial exclusion. Together, these contributions provide the first comprehensive picture of how bistability, geometry, and diffusion interact to shape coexistence in predator-prey systems with exclusion zones.
2. Existence of non-trivial equilibria
This section and the following one are devoted to the proofs of Theorems 1.3 and 1.4, about the existence of non-trivial equilibria of the predator-prey system with exclusion. That is, stationary solutions that are non-constant with a non-zero prey population. There are indeed several constant as well as non-constant solutions with $v\equiv 0$. In order to avoid these and ascertain that we have identified non-trivial solutions, we employ a construction that rests on a topological degree argument.
We first detail the proof in the one-dimensional case, that is for the system (2). It is carried in several steps.
Step 1. Formulation as a fixed point equation.
We first observe that the second equation (or vv–equation) of the system is linear in $v$. Thus, a positive solution $v$ of this equation can be interpreted as a principal eigenfunction of an operator with a potential term. Since such a solution is unique up to multiplication by a constant, this leads us to formulate the problem for a pair $(u,s)$ where $s$ is a scalar.
More precisely, first, for $u\in C^{1}[0,L]$ we define $\lambda=\lambda[u]$ and $\phi=\phi[u]$ as the principal eigenvalue and eigenfunction of the problem:
[…]
The last requirement is a normalization condition. Thus, the second equation simply reads $\lambda[u]=0$ with $v$ some positive multiple of $\phi$, that is $v=s\phi[u]$ for some $s>0$.
We can now formulate the problem with unknowns $(u,s)$. Consider the function space
$\mathcal{E}:=C^{1}[0,L]\times\mathbb{R}.$
Let $K$ be a positive constant chosen appropriately large so that $z\mapsto f(z)+Kz-\beta Sz$ is increasing over the interval $z\in[0,1]$ where the choice of the constant $S$ is discussed below. For $(u,s)\in\mathcal{E}$, we define $w:=\mathcal{F}(u,s)$ as the unique solution of
[…]
$\begin{cases}&u=\mathcal{F}(u,s),\\ &\lambda[u]=0.\end{cases}$ (7)
$\mathcal{G}(u,s):=(\mathcal{F}(u,s),s-\lambda[u]),$
Lastly, since we are looking for positive solutions, we restrict the pair $(u,s)$ to the open set (in $\mathcal{E}$) defined by $0<u<1$ in $(0,L)$ and $0<s<S$ for some positive constant $S$ appropriately large to be determined later. Actually, we are going to look for a solution $(u,s)$ such that $u>\zeta$ on $(a,L)$ where $\zeta$ comes from Proposition 1.2. Let us extend $\zeta$ to $[0,L]$ by setting $\zeta(x)=0$ for all $x\in(0,a)$. Thus, we define $\mathcal{D}\subset\mathcal{E}$ by
$\mathcal{D}=D\times(0,S),\quad\text{with}\quad D:=\{u\in C^{1}[0,L]\,;\;\zeta<u<1\text{ in }(0,L)\}.$
Step 2. Topological degree
We will show the existence of a solution using a topological degree argument.
Let us set
It is well known that $H$ is a compact perturbation of the identity in $\mathcal{E}$, that is, $\mathcal{F}$ is a compact mapping $\mathcal{E}\rightarrow C^{1}[0,L]$ and $\lambda$ is continuous and bounded on bounded sets in the space $C^{1}[0,L]$. We will show that the Leray-Schauder topological degree
$d(H,\mathcal{D},(0,0))$
of the mapping $H$ in the open set $\mathcal{D}$ with target point $(0,0)$ is nonzero. We first need to show that the degree is well defined. This amounts to showing the following.
Lemma 2.1.
The mapping $H$ does not hit the target value $(0,0)$ on the boundary of $\mathcal{D}$.
For the proof, we require the observations in the following two lemmas.
Lemma 2.2.
When LaL-a is sufficiently large, a solution $u$ of
$\begin{cases}d_{u}u_{xx}+f(u)=0\quad\text{in}\quad(0,L),\\ u_{x}(0)=u_{x}(L)=0\end{cases}$ (8)
that satisfies $u\geq\zeta$ on $(a,L)$ is none other than the trivial solution $u\equiv 1$.
Proof.
Recall that $\zeta$ comes from Proposition 1.2, and is the maximal positive solution of (3), which exists provided LaL-a is sufficiently large. Fix $a>0$ and assume $L$ is sufficiently large that such a solution $\zeta$ of (3) exists. From the equation, we get:
$\int_{0}^{\zeta(L)}f(s)ds\,=\,\frac{d_{u}}{2}\zeta^{\prime}(a)^{2}>0.$
This entails that $\zeta(L)>\rho$ where $0<\rho<1$ is defined by $\int_{0}^{\rho}f(s)ds=0$. From equation (8), we see that
$\int_{u(0)}^{u(L)}f(s)ds=0.$
Since $u(L)\geq\zeta(L)>\rho$, this implies that $u(0)=u(L)$ from which it is straightforward to obtain that $u$ is constant and thus that $u\equiv 1$. ∎
The following observation will be useful.
Lemma 2.3.
There exists $\delta>0$ such that for all $u$ with $0\leq u\leq 1$, the eigenfunction defined by (5) satisfies
$0<\delta\leq\phi[u](x)\leq 1\quad\text{for all}\quad x\in[0,a]\quad\text{and}\quad u\in\overline{D}.$
Proof.
We know that $\lambda[u]$ is bounded, and, to be more precise, that $\gamma-\alpha\leq\lambda[u]\leq\gamma$ for $0\leq u\leq 1$. Therefore, from the eigenfunction equation (5), and by the normalization condition $\sup_{x\in[0,a]}\phi[u](x)=1$ it follows that $\phi[u]$ is bounded from above and away from 0. Indeed, otherwise, one could find a sequence of functions $u_{j}$ with $0\leq u_{j}\leq 1$, and corresponding eigenvalues and eigenfunctions $\lambda_{j}=\lambda[u_{j}],\phi_{j}=\phi[u_{j}]$ solutions of the eigenvalue problem (5) and such that $\min_{[0,a]}\phi_{j}\searrow 0$ as $j\nearrow\infty$. We can strike out subsequences (still denoted with index $j$) such that
$\lambda_{j}\to\lambda,\quad u_{j}\rightharpoonup u\quad\text{in}\quad L^{p},\quad\phi_{j}\rightharpoonup\phi\quad\text{in}\quad W^{2,p},\quad\text{for all}\quad p>1.$
By classical embedding, we know that $\phi_{j}\to\phi$ in $C^{1}[0,a]$. Therefore, $\phi\in W^{2,\infty}$ satisfies
$\phi_{xx}=q(x)\phi,\quad\phi\geq 0,\quad\text{in}\quad(0,a),\quad\text{and}\quad\phi_{x}(0)=\phi_{x}(a)=0.$
for some bounded function $q\in L^{\infty}(0,a)$. Moreover,there must exist a point $x_{0}\in[0,a]$ where $\phi$ vanishes. In view of the equation above, this implies $\phi\equiv 0$, which contradicts the fact that owing to the $C^{1}$ convergence, we have $\max\phi=1$.
Therefore, there is a $\delta>0$ such that
$0<\delta\leq\phi[u](x)\leq 1\quad\text{for all}\quad x\in[0,a]\quad\text{and}\quad u\in\overline{D}.$
The next result provides an a priori bound on $s$ for the solutions of system (7).
Lemma 2.4.
For any $\gamma_{0}$ with $0<\gamma_{0}<\gamma$, there exists $S=S(\gamma_{0})>0$ (sufficiently large) such that any solution $(u,s)$ of equation (6) with $s\geq S$ and $u\in\overline{D}$ must satisfy $\lambda[u]\geq\gamma_{0}$.
Proof.
We know that $f(u)\leq\sigma_{0}u$ for all $u\in\overline{D}$ for some constant $\sigma_{0}>0$. Therefore, if $(u,s)\in\mathcal{D}$ is a solution with $s\geq S$, using the previous lemma, we have
$d_{u}u_{xx}\leq[\sigma_{0}-\beta S\delta]u\quad\text{on}\quad[0,a].$
Hence, for any $\kappa>0$ one may choose $S$ sufficiently large so that
$u_{xx}\geq\kappa^{2}u\quad\text{on}\quad[0,a].$ (9)
With this choice of $S$ the previous inequality yields
$u(x)\geq u(0)\cosh{{\kappa}x}.$
In fact, for any $x\in[0,a]$, we have
$u(y)\geq u(x)\cosh{\kappa(y-x)}.$
Using the fact that $u\leq 1$ we infer that, given any $\eta>0$ arbitrarily small, we can choose $S$ sufficiently large and consequently $\kappa$ sufficiently large so that the previous inequalities imply
$u(x)\leq\eta\quad\text{for all}\quad x\in(0,a-\eta).$ (10)
Let $\lambda[u],\phi[u]$ be the eigen-pair defined by equation (5). In view of Lemma 2.3, for any $0<\epsilon<1$, one can find $\eta$ (independent of $u$) such that
$\int_{0}^{a-\eta}\phi^{2}\;\geq(1-\epsilon)\int_{0}^{a}\phi^{2},\quad\text{and}\quad\int_{a-\eta}^{a}\phi^{2}\leq\epsilon\int_{0}^{a}\phi^{2}.$
From the relation:
$\lambda[u]=\frac{\int_{0}^{a}d_{u}\phi^{\prime 2}+q(x)\phi^{2}}{\int_{0}^{a}\phi^{2}}$
where $q(x):=(\gamma-\alpha u)$, we then infer
$\lambda[u]\geq(\gamma-\alpha\eta)(1-\epsilon)-(\alpha-\gamma)\epsilon.$
The lower bound on $\lambda[u]$ of the lemma then follows by choosing $\epsilon$ and $\eta$ sufficiently small. ∎
We can now complete the proof of Lemma 2.1. Fix some $\gamma_{0}\in(0,\gamma)$ and for the definition of $\mathcal{D}$ choose $S=S(\gamma_{0})$ as in the previous lemma. We claim that $H\not=(0,0)$ on the boundary of the set $\mathcal{D}$.
Let $(u,s)\in\partial\mathcal{D}$ be a solution of system (7). There are three possibilities.
(i) $s=0,u\in\overline{D}$. Then, owing to Lemma 2.2, $u\equiv 1$. But we know that $\lambda[1]=\gamma-\alpha<0$ and $(u\equiv 1,s=0)$ is not a solution of (7).
(ii) $s=S,u\in\overline{D}$. By our choice of $S$, in view of Lemma 2.4, we know that $\lambda[u]\geq\gamma_{0}>0$ (where we have chosen some value $0<\gamma_{0}<\gamma$). Thus, $(u,S)$ cannot be a solution of (7) either.
(iii) $0\leq s\leq S,u\in\partial{D}$. By definition of $D$, the solution $u$ satisfies $\zeta\leq u\leq 1$ on $[0,L]$. We know by step (i) that $s>0$ which entails that $u\not\equiv 1$ and we also know (by definition of $D$) that $u\not\equiv 0$. We then infer from the strong maximum principle that $0<u<1$ on the closed interval $[0,L]$. For the inequality $u<1$ we use the fact that since $s>0$, $u$ is a strict sub-solution of the equation $d_{u}w_{xx}+f(w)=0$ on $(0,L)$ and we also use the strong maximum principle at the end points of the interval. Lastly, $u$ and $\zeta$ are solutions of that same equation on the interval $(a,L)$. Since $u\geq\zeta$ on $(a,L)$, with $u_{x}(L)=\zeta_{x}(L)=0$, and $u(a)>0=\zeta(a)$, the strong maximum principle also shows that $u>\zeta$ on $[a,L]$. Therefore, the function $u$ is in the interior of the set $D$ and this case is ruled out as well. ∎
Step 3. Computation of the topological degree
The existence of a solution of (7) with $u\geq\zeta$ on $(a,L)$ is is a consequence of the fact that the topological degree $d(H,\mathcal{D},(0,0))$ is not zero. We now compute this degree.
Proposition 2.5.
$d(H,\mathcal{D},(0,0))=1.$
Proof.
Let $\sigma$ be a $C^{1}$ function defined on $\mathbb{R}^{+}$ such that $\sigma^{\prime}<0$, on $(0,S)$, $\sigma(0)=1$ and $\sigma(s)=0$ for all $s\geq S$ where $S$ comes from the definition of $\mathcal{D}$. We consider a family of deformation of $H$ parametrized by $\tau\in[0,1]$ defined by
$H_{\tau}(u,s)\,:=\,\left(\,u-\mathcal{F}(u,s)\,,\,\lambda[(1-\tau)u+\tau\sigma(s)]\,\right).$
Thus, we have $H_{0}=H$. Let us denote $H_{\tau}(u,s)=(\tilde{H}(u,s),h_{\tau}(u,s))$ where
$\tilde{H}(u,s):=u-\mathcal{F}(u,s)\,\quad\text{and}\quad h_{\tau}(u,s)=\lambda[(1-\tau)u+\tau\sigma(s)].$
We need to extend Lemma 2.1 to the homotopy $H_{\tau}$.
Lemma 2.6.
$H_{\tau}(u,s)\neq(0,0)$ on $\partial\mathcal{D}.$
Proof.
Let $(u,s)\in\partial\mathcal{D}$ satisfy $H_{\tau}(u,s)=(0,0)$ for some $\tau\in[0,1]$. We rule this out by examining the same three cases as before.
(i) $s=0,u\in\overline{D}$. Then, since $\tilde{H}$ does not depend on $\tau$, as before Lemma 2.2 implies that $u\equiv 1$. Then, $h_{\tau}(u,s)=h_{\tau}(1,0)=\lambda[1]=\gamma-\alpha<0$ and this case cannot happen.
(ii) $s=S,u\in\overline{D}$. We know that $h_{\tau}(u,S)=\lambda[(1-\tau)u]\geq\lambda[u].$ Lemma 2.4 again guarantees that $h_{\tau}(u,S)>0$. Note that, indeed, the first component of $H_{\tau}$, i.e. $\tilde{H}$ is independent of $\tau$ and is the same as that of $H$. Thus, $(u,S)$ cannot be a solution of (7) either.
(iii) $0\leq s\leq S,u\in\partial{D}$. By step (i) we know that $s>0$. Then we only use the first equation $\tilde{H}(u,s)=0$ which is the same as the one we discussed in part (iii) of the proof of Lemma 2.1 and the proof is unchanged. ∎
We can now conclude the proof of Proposition 2.5. The homotopy invariance of the Leray-Schauder degree implies that
$d(\,H,\mathcal{D},(0,0)\,)\,=\,d(\,H_{1},\mathcal{D},(0,0)\,).$
Now, $H_{1}$ is in triangular form, because its second component, $h_{1}(u,s)$ does not depend on $u$. Indeed, we have
$h_{1}(u,s)=\lambda[\sigma(s)]=\gamma-\alpha\sigma(s).$
Since $\gamma<\alpha$ and $\sigma^{\prime}<0$ on $(0,S)$, this last equation has a unique solution $s=s_{1}\in(0,S)$. This shows that we can separate the operator in its two components, that is “diagonalize it”, to obtain [3]:
We know that $d(h_{1},(0,S),0)=1$.
Let us compute the first factor in the product in the right-hand side. To this end, we observe the following
Lemma 2.7.
Let $K$ be such that the function $z\mapsto f(z)+Kz-\beta Sz$ is increasing on $[0,1]$. Then, $\mathcal{F}(\cdot,s_{1})$ maps $D$ into its interior.
Proof.
This motivates the choice of $K$ in the definition of the operator $\mathcal{F}$. Let $\zeta\leq u\leq 1$ on $(0,L)$. Recall that $\zeta$ is extended by 0 over $(0,a)$ and that $w=\mathcal{F}(u,s_{1})$ is defined as the solution of the problem:
[…]
By the choice of $K$, we know that for such a $u$ we have
$\begin{cases}f(u)+Ku-\beta s_{1}\phi[u]u\geq 0&\quad\text{for}\quad x\in(0,a),\\ f(u)+Ku\geq f(\zeta)+K\zeta&\quad\text{on}\quad[a,L],\\ f(u)+Ku-\beta s_{1}\phi[u]u<f(1)+K=K&\quad\text{for}\quad x\in(0,L).\end{cases}$
From the last inequality, we get
$\begin{cases}-d_{u}(w-1)_{xx}+K(w-1)<0\quad\text{in}\quad(0,L),\\ (w-1)_{x}(0)=(w-1)_{x}(L)=0,\end{cases}$
whence, by the maximum principle, $w<1$ on $[0,L]$. Likewise, since $d_{u}w_{xx}+Kw\geq 0$ on $[0,L]$ and $d_{u}(w-\zeta)_{xx}+K(w-\zeta)\geq 0$ on $(a,L)$, we get $w\geq 0$ on $[0,L]$ and $w\geq\zeta$ on $[a,L]$. From the fact that $\zeta$ extended by 0 on $[0,a]$ is a strict sub-solution of the equation, we derive that these inequalities are strict and therefore that $w$ is in the interior of $D$. ∎
$d(\,H,\mathcal{D},(0,0)\,)\,=\,1.$
which establishes the claim of the proposition 1.2.
We note the contrast between Theorem 1.3 and Proposition 1.1 in which we show that if $\gamma>\alpha$, then the predators go extinct, so there is no coexistence equilibrium. The case when the predation zone is contained in the interior of the predation zone in 1D is a special case of the existence theorem for arbitrary dimensions: Theorem 1.4, proved in Section 3.
3. Coexistence equilibrium in higher dimensions
This section is devoted to the proof of Theorem 1.4. It relies on the same type of fixed point formulation and topological degree argument as in the one-dimensional case of the previous section but it requires several new PDE ingredients.
We note by $R:=\rho+R_{0}$ so that $A\subset B_{\rho}\subset B_{R}\subset\Omega$. Without loss of generality, we assume that each of these balls is centered at the origin. We want to show that (4) admits a coexistence solution $(u,v)$ provided $R_{0}$ exceeds some threshold value. The proof will be carried out in a sequence of steps.
@@ 3. Coexistence equilibrium in higher dimensions
We now apply the sliding method as in [7] to show that as we shift the subsolution $V(\cdot-a)$ by moving around $a$, it keeps remaining below $u$ in $B_{R}$. More precisely, we claim that $u(b)>m$ for an arbitrary point $b\in B_{R}$. To this end, we extend $V$ by 0 outside $B_{\sigma}(0)$, let $a_{t}=(1-t)a+tb$ and set:
$t^{*}=\max\;\{\;\overline{t}\in[0,1]\,;\,u(x)\geq V(x-a_{t})\quad\text{for all}\quad x\in B_{R},\;t\in[0,\overline{t}]\;\}.$
We know that $t^{*}\geq 0$. Note that the segment from $a$ to $b$ lies in the ball $B_{R}$ but may cross the ball $B_{\rho}$, and the sliding ball $B_{\sigma}(a_{t})$ may have portions outside $B_{R}$. Nonetheless, we claim that $t^{*}=1$. For if not, for some $t^{*}\in[0,1)$, we would have:
$u(x_{0})-V(x_{0}-a_{t^{*}})=\min\,\{\,u(x)-V(x-a_{t^{*}})\,;\,x\in B_{R}\,\}\,=\,0,$
for some $x_{0}\in\overline{B_{R}}$. On $\partial B_{R}$, by construction we have $u\geq 1-\eta>\max V$. We also know that $u(x_{0})>0$. Therefore, $x_{0}$ must be in $B_{R}\cap B_{\sigma}(a_{t^{*}})$ and it is thus an interior minimum. Since both $u$ and $V$ satisfy the same semi-linear elliptic equation, the strong maximum principle (or positivity property) shows that $u\equiv V$ which is impossible.
Hence, it must be the case that $t^{*}=1$, whence it follows that $u(b)\geq m=\max V$. (Actually, the proof shows that $u(b)>m$.) Since $b$ is arbitrary, we have proved that $\min_{B_{R}}u\geq m$. But we also know that $u\geq 1-\eta>m$ in $\Omega\setminus B_{R}$. Together these inequalities show that $u\geq m$ in all of $\overline{\Omega}$. This entails that $\Delta u\geq 0$ in $\Omega$. The function $u$ reaches its minimum at some point $x_{0}\in\overline{\Omega}$. By the strong maximum principle (in the interior or at the boundary) we derive that $u$ is constant and since $u\geq m>\theta$, this can only happen when $u\equiv 1$. The proof of the Lemma is thereby complete. ∎
@@
As in the one-dimensional case, we formulate the system as a fixed point equation. For $u\in{\mathrm{C}}^{1}(\overline{\Omega})$ we define $\lambda=\lambda[u]$ and $\phi=\phi[u]$ as the principal eigenvalue and eigenfunction of the problem:
@@
$H_{\tau}(u,s)\,:=\,\left(\,u-\mathcal{F}(u,s)\,,\,\lambda[(1-t)u+t\sigma(s)]\,\right),\quad t\in[0,1].$
Thus, we have $H_{0}=H$. The mapping $H_{\tau}$ is a compact perturbation of the identity. In fact, more precisely, it can be written in the form $H_{\tau}=Id_{|\mathcal{E}}-\mathfrak{F}_{t}$ with $\mathfrak{F}$ a continuous mapping from $[0,1]$ into the space of compact operators on $\overline{D}$.
@@
Case 3. $u(x_{0})=1$ for some $x_{0}\in\overline{\Omega}$. We notice that $u$ satisfies a linear equation $d_{u}\Delta u=f(u)-q(x)u$ for some $L^{\infty}$ function $q\geq 0$, while $1$ is a supersolution of the same equation with $1\geq u$. Moreover, by the Agmon-Douglis-Nirenberg estimate, $u$ is of class $W^{2,p}(\Omega)$ for all $p<\infty$, and therefore, if $x_{0}\in\Omega$, by the strong maximum principle we get $u\equiv 1$ in $\Omega$. Now, $u$ is of class $\mathrm{C}^{2}$ near the boundary of $\Omega$. Hence, from the strong maximum principle at the boundary we infer $u\equiv 1$ as well if $x_{0}\in\partial\Omega$. But for $u\equiv 1$ to solve the equation, we must have $s=0$. Then, we get $h_{\tau}(u,s):=\lambda[(1-t)u+t\sigma(s)]=\lambda[1-t+t\sigma(0)]=\lambda[1]=\gamma-\alpha<0$. Hence, $u\equiv 1$ cannot be a solution and we have shown that $u<1$ on $\overline{\Omega}$.
@@
The solution $w=w_{\kappa}$ of equation (22) for $\kappa>0$ satisfies $0<w_{\kappa}<1$ and $w_{\kappa}\searrow 0$ as $\kappa\nearrow\infty$, uniformly on compact subsets of $A$.
@@
Hence, $\overline{z}$ is a supersolution of equation (23) and thus $z(0)\leq\overline{z}(0)=e^{-\xi\delta^{2}}$. It follows that, given $\delta,\epsilon>0$, there is a $\kappa_{0}$ sufficiently large such that, for any $\kappa\geq\kappa_{0}$, we have $z(0)\leq\epsilon$.
Now, considering the set $A$ and the equation (22), for any point $x_{0}\in A$ whose distance to the boundary is larger than $\delta$, we can compare $w$ and $z(\cdot-x_{0})$ which implies $w(x_{0})<e^{-\xi\delta^{2}}$. We have thus proved that for any $\delta>0$, the solution $w$ converges uniformly to 0 on the set $\{x\in A\,;\,\mathrm{dist}(x,\partial A)\geq\delta\}$. ∎
@@
Comparison with (22) yields $u\leq w$ in $A$. Therefore, given some $\epsilon>0$ and $\delta>0$, if $S$ is sufficiently large, we get $u\leq\epsilon$ in the set $A_{\delta}:=\{x\in A\,;\,\mathrm{dist}(x,\partial A)\geq\delta\}$ while $u\leq 1$ in the complement.
@@
Splitting the second integral over the set $A_{\delta}$ and its complement $N_{\delta}$, which is the $\delta$-neighborhood of the boundary in $A$, and using Lemma 3.8 yield:
$\frac{\int_{A}u\phi^{2}}{\int_{A}\phi^{2}}\leq\frac{1}{\eta}\frac{|N_{\delta}|}{|A|}+\epsilon.$ (26)
Given $0<\epsilon<(\gamma-\gamma_{0})/2$, we choose $\delta>0$ small enough so that $\frac{1}{\eta}\frac{|N_{\delta}|}{|A|}<\epsilon$. From this choice of $\epsilon,\delta>0$, we get a value of $S$ such that for any $s\geq S$, from (25) and (26) we infer $\lambda-\gamma\geq 2\epsilon$ which yields $\lambda[u]\geq\gamma_{0}$. This completes the proof of Lemma 3.7 which allows us to rule out case 5. ∎
@@
The existence of a solution of (7) with $u\geq\zeta$ on $(a,L)$ is is a consequence of the fact that the topological degree $d(H,\mathcal{D},(0,0))$ is not zero. We now compute this degree.

### v2

@@ 1.4 Related literature
Specifically, our paper addresses the gaps in the literature through three principal advances. First, we establish the existence of coexistence equilibria without appealing to bifurcation from semi-trivial states. Using a novel topological-degree argument, we prove the existence of nontrivial steady states in parameter regimes that may be remote from classical bifurcation points. This yields a more global and geometrically robust understanding of coexistence, complementing and extending the local bifurcation analyses of, e.g., [13, 11, 35, 56]. Second, we develop a rigorous thin-limit analysis showing how coexistence persists and transforms as the predator region becomes vanishingly thin. This asymptotic framework connects the fully coupled PDE model to lower-dimensional effective equations, revealing new behaviors not captured by prior protection-zone studies. Finally, our numerical investigations suggest the occurrence of a variety of nonlinear phenomena including multiple coexistence branches, turning-point bifurcations, Hopf bifurcations, and transitions between extinction and persistence regimes. These phenomena highlight the rich dynamical landscape made possible by the coupling of Allee-type prey dynamics and spatial exclusion. Together, these contributions provide the first comprehensive picture of how bistability, geometry, and diffusion interact to shape coexistence in predator–prey systems with exclusion zones.
2 Coexistence equilibrium in higher dimensions
This section is devoted to the proof of Theorem 1.2. It relies on a fixed point formulation and topological degree argument. We denote by $R:=\rho+R_{0}$ so that $A\subset B_{\rho}\subset B_{R}\subset\Omega$. Without loss of generality, we assume that each of these balls is centered at the origin. We want to show that (2) admits a coexistence solution $(u,v)$ provided $R_{0}$ exceeds some threshold value. The proof will be carried out in a sequence of steps.
@@ 2 Coexistence equilibrium in higher dimensions
We now apply the sliding method as in [6] to show that as we shift the subsolution $V(\cdot-a)$ by moving around $a$, it keeps remaining below $u$ in $B_{R}$. More precisely, we claim that $u(b)>m$ for an arbitrary point $b\in B_{R}$. To this end, we extend $V$ by 0 outside $B_{\sigma}(0)$, let $a_{s}=(1-s)a+sb$ and set:
$s^{*}=\max\;\{\;\overline{s}\in[0,1]\,;\,u(x)\geq V(x-a_{s})\quad\text{for all}\quad x\in B_{R},\;s\in[0,\overline{t}]\;\}.$
We know that $t^{*}\geq 0$. Note that the segment from $a$ to $b$ lies in the ball $B_{R}$ but may cross the ball $B_{\rho}$, and the sliding ball $B_{\sigma}(a_{s})$ may have portions outside $B_{R}$. Nonetheless, we claim that $s^{*}=1$. For if not, for some $s^{*}\in[0,1)$, we would have:
$u(x_{0})-V(x_{0}-a_{s^{*}})=\min\,\{\,u(x)-V(x-a_{s^{*}})\,;\,x\in B_{R}\,\}\,=\,0,$
for some $x_{0}\in\overline{B_{R}}$. On $\partial B_{R}$, by construction we have $u\geq 1-\eta>\max V$. We also know that $u(x_{0})>0$. Therefore, $x_{0}$ must be in $B_{R}\cap B_{\sigma}(a_{s^{*}})$ and it is thus an interior minimum. Since both $u$ and $V$ satisfy the same semi-linear elliptic equation, the strong maximum principle (or positivity property) shows that $u\equiv V$ which is impossible.
Hence, it must be the case that $s^{*}=1$, whence it follows that $u(b)\geq m=\max V$. (Actually, the proof shows that $u(b)>m$.) Since $b$ is arbitrary, we have proved that $\min_{B_{R}}u\geq m$. But we also know that $u\geq 1-\eta>m$ in $\Omega\setminus B_{R}$. Together these inequalities show that $u\geq m$ in all of $\overline{\Omega}$. This entails that $\Delta u\geq 0$ in $\Omega$. The function $u$ reaches its minimum at some point $x_{0}\in\overline{\Omega}$. By the strong maximum principle (in the interior or at the boundary) we derive that $u$ is constant and since $u\geq m>\theta$, this can only happen when $u\equiv 1$. The proof of the Lemma is thereby complete. ∎
@@
We formulate the system as a fixed point equation. For $u\in{\mathrm{C}}^{1}(\overline{\Omega})$ we define $\lambda=\lambda[u]$ and $\phi=\phi[u]$ as the principal eigenvalue and eigenfunction of the problem:
@@
$H_{\tau}(u,s)\,:=\,\left(\,u-\mathcal{F}(u,s)\,,\,\lambda[(1-\tau)u+\tau\sigma(s)]\,\right),\quad\tau\in[0,1].$
Thus, we have $H_{0}=H$. The mapping $H_{\tau}$ is a compact perturbation of the identity. In fact, more precisely, it can be written in the form $H_{\tau}=\mathrm{Id}_{|\mathcal{E}}-\mathfrak{F}_{\tau}$ with $\mathfrak{F}$ a continuous mapping from $[0,1]$ into the space of compact operators on $\overline{D}$.
@@
Case 3. $u(x_{0})=1$ for some $x_{0}\in\overline{\Omega}$. We notice that $u$ satisfies a linear equation $d_{u}\Delta u=f(u)-q(x)u$ for some $L^{\infty}$ function $q\geq 0$, while 11 is a supersolution of the same equation with $1\geq u$. Moreover, by the Agmon-Douglis-Nirenberg estimate, $u$ is of class $W^{2,p}(\Omega)$ for all $p<\infty$, and therefore, if $x_{0}\in\Omega$, by the strong maximum principle we get $u\equiv 1$ in $\Omega$. Now, $u$ is of class $\mathrm{C}^{2}$ near the boundary of $\Omega$. Hence, from the strong maximum principle at the boundary we infer $u\equiv 1$ as well if $x_{0}\in\partial\Omega$. But for $u\equiv 1$ to solve the equation, we must have $s=0$. Then, we get $h_{\tau}(u,s):=\lambda[(1-\tau)u+\tau\sigma(s)]=\lambda[1-\tau+\tau\sigma(0)]=\lambda[1]=\gamma-\alpha<0$. Hence, $u\equiv 1$ cannot be a solution and we have shown that $u<1$ on $\overline{\Omega}$.
@@
$0<\eta\leq\phi[u](x)\leq 1\quad\text{for all}\quad x\in A\quad\text{and}\quad u\in\overline{D}.$
@@
The solution $w=w_{\kappa}$ of equation (13) for $\kappa>0$ satisfies $0<w_{\kappa}<1$ and $w_{\kappa}\searrow 0$ as $\kappa\nearrow\infty$, uniformly on compact subsets of AA.
@@
Hence, $\overline{z}$ is a supersolution of equation (14) and thus $z(0)\leq\overline{z}(0)=e^{-\xi\delta^{2}}$. It follows that, given $\delta,\varepsilon>0$, there is a $\kappa_{0}$ sufficiently large such that, for any $\kappa\geq\kappa_{0}$, we have $z(0)\leq\varepsilon$.
Now, considering the set AA and the equation (13), for any point $x_{0}\in A$ whose distance to the boundary is larger than $\delta$, we can compare $w$ and $z(\cdot-x_{0})$ which implies $w(x_{0})<e^{-\xi\delta^{2}}$. We have thus proved that for any $\delta>0$, the solution $w$ converges uniformly to 0 on the set $\{x\in A\,;\,\mathrm{dist}(x,\partial A)\geq\delta\}$. ∎
@@
Comparison with (13) yields $u\leq w$ in AA. Therefore, given some $\varepsilon>0$ and $\delta>0$, if $S$ is sufficiently large, we get $u\leq\varepsilon$ in the set $A_{\delta}:=\{x\in A\,;\,\mathrm{dist}(x,\partial A)\geq\delta\}$ while $u\leq 1$ in the complement.
@@
Splitting the second integral over the set $A_{\delta}$ and its complement $N_{\delta}$, which is the $\delta$-neighborhood of the boundary in AA, and using Lemma 2.8 yield:
$\frac{\int_{A}u\phi^{2}}{\int_{A}\phi^{2}}\leq\frac{1}{\eta}\frac{|N_{\delta}|}{|A|}+\varepsilon.$ (17)
Given $0<\varepsilon<(\gamma-\gamma_{0})/2$, we choose $\delta>0$ small enough so that $\frac{1}{\eta}\frac{|N_{\delta}|}{|A|}<\varepsilon$. From this choice of $\varepsilon,\delta>0$, we get a value of $S$ such that for any $s\geq S$, from (16) and (17) we infer $\lambda-\gamma\geq 2\varepsilon$ which yields $\lambda[u]\geq\gamma_{0}$. This completes the proof of Lemma 2.7 which allows us to rule out case 5. ∎
@@
The existence of a solution of (2) with $u\geq\zeta$ on $\Omega\setminus A$ is a consequence of the fact that the topological degree $d(H,\mathcal{D},(0,0))$ is not zero. We now compute this degree.